8.02 - ESG Independent Study Spring 2005

Syllalbus

There are sixteen units in ESG’s 8.02 - Independent Study sequence. All sixteen units must be passed in order to complete the subject.

Unit 1  Electrostatics
Unit 2  Gauss’ Law
Unit 3  Electric Potential
Unit 4  Work and Energy
Unit 5  Conductors
Unit 6  Capacitance and Dielectrics
Unit 7  Current and Resistance
Unit 8  Electromotive Force and Circuits
Unit 9  The Magnetic Field
Unit 10  Ampère’s Law
Unit 11  Faraday’s Law
Unit 12  Inductance
Unit 13  Magnetic Properties of Matter
Unit 14  AC Circuits
Unit 15  Maxwell’s Equations
Unit 16  Electromagnetic Waves and Light

Each of the sixteen units has a study guide with suggested readings and problems from one of several texts (see the Introduction for details on which texts are currently in use for 8.02IS). Units are passed by satisfactory completion of a unit test. These tests are primarily written, but may involve oral explanation of simple E&M phenomena.